OpenStudy (anonymous):

Simplify With All Positive Exponents : (u^-2/ w^3)(-2u^2v^3w)^-3

OpenStudy (anonymous):

\[\frac{u^{-2}}{ w^3}\times( -2u^2v^3w)^{-3}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{1}{u^2w^3}\times(-2)^{-3}u^{-6}v^{-9}w^{-3}\]

OpenStudy (anonymous):

\[=\frac{1}{u^2w^3(-8)u^6v^9w^3}\]

OpenStudy (anonymous):

\[=\frac{1}{-8u^8v^9w^6}\]